Please help me with this problem:
The cup company charges $4 per mug plus an additional fee of $30. Mug Rus only charges $4.50 per mug.
Write and solve and equation to find how many mugs Manuela can have made for the 2 companies prices to be the same.
thanks

Answer by ankor@dixie-net.com(22740)

let m = no. of mugs
write a cost equation for each company
The cup company charges $4 per mug plus an additional fee of $30.
C = 4m + 30
and
Mug Rus only charges $4.50 per mug.
C = 4.50m
:
Write and solve an equation to find how many mugs Manuela can have made for the 2 companies prices to be the same.
they want the two costs to be same therefore:
4.50m = 4m + 30
4.50m - 4m = 30
.50m = 30
m = 30/.5
m = 60 mugs, the cost will be the same
:
:
Confirm this; find the actual cost for each
C = 4(60) + 30
C = 240 + 30
C = $270
and
C = 4.50(60)
C = $270